OpenStudy (anonymous):

Elastic Collisions: "A charging bull elephant with a mass of 5400kg comes directly toward you with a speed of 4.30 m/s. You toss a 0.150 kg ball at the elephant with a speed of 8.11 m/s. " (a) When the ball bounces back toward you, what is its speed? (b) How do you account for the fact that the ball's kinetic energy has increased? **The answer in the back of the book for (a) is 17 m/s.**

OpenStudy (anonymous):

I got 16.7 m/s perhaps only approximate answer. Started with conservation of momentum M v1 - m v2 = M v1' + m v2' where primes mean after collision Then used conservation of energy 1/2 M v1^2 + 1/2 m v2^2 = 1/2 M v1'^2 + 1/2 m v2'^2 The put the Ms on left side of each equation and the ms on the right. Divided the momentum equation into the energy equation, noting (a^2-b^2)=(a+b)(a-b) Assumed M so much larger than m so that v1+v1' = 2 v1, got 2(4.30) = v2' - 8.11 v2' = 16.7. The extra energy comes from the motion of the beast, slightly slowed.

OpenStudy (anonymous):

NOTE! This way definitely works, but it's really, really ugly :P You can do it without assuming infinite mass of the elephant (although granted it doesn't change the answer much) Starting with your conservation of momentum \[p_i=p_f\] \[Mv_1 - mv_2 = Mv_1' + mv_2'\] \[\qquad \qquad \qquad A= (Mv_1-mv_2) \quad \leftarrow \text{do this to make our lives easier!}\] \[ A = Mv_1' - mv_2'\] \[Mv_1' = A-mv_2'\] \[(Mv_1')^2 = (A-mv_2')^2\] \[Mv_1^{' \ 2}=\frac{(A-mv_2')^2}{M}\] Then in your conservation of energy (just drop all of the halves off of the bat since every term has them) \[Mv_1^2 + mv_2^2 = Mv_1^{' \ 2} + mv_2^{' \ 2}\] sub in our M(v1')^2 from above \[Mv_1^2 + mv_2^2 = \frac{(A-mv_2')^2}{M} + mv_2^{' \ 2}\] So now solve or v2' - the velocity of the ball \[ Mm(v_2')^2 +A^2 - 2Amv_2' + (mv_2')^2= MMv_1^2 + Mmv_2^2\] ick!! \[Mmv_2^{' \ 2}+(mv_2')^2-2Amv_2'+A^2-(MMv_1^2 + Mmv_2^2)=0\] \[v_2^{' \ 2}(Mm+m^2)-2Amv_2'+A^2-(MMv_1^2 + Mmv_2^2)=0\] triple ick. We can see that it's a quadratic on v_2' however, so \[ (v_2')^2 \Big( Mm+m^2 \Big) + v_2' \Big( -2Am \Big) + \Big( A^2 - (MMv_1^2 + Mmv_2^2) \Big) = 0\] to use in \[v_2' = \frac{-b±\sqrt{b^2-4ac}}{2a}\] quadruple ick \[a = Mm+m^2 = ((5400kg)(.150kg)+(.150kg)^2) \approx 795kg^2 \] \[b = (-2Am) = 2(Mv_1' + mv_2')(m) \\ = 2((5400kg)(4.3m/s)-(.15kg)(8.11m/.s))(.15kg) \\ b \approx -6966 kg^2 m/s\] \[c = A^2-(MMv_1^2 + Mmv_2^2) \\ =((5400kg)(4.3m/s)-(.15kg)(8.11m/.s))^2-(5400kg)^2(4.30m/s)^2 \\ \qquad - (5400kg)(.15kg)(8.11m/s)^2 \\ c \approx -109768 kg^2 m^2/s^2\] Just yuck! then we get \[v_2' = \frac{-b±\sqrt{b^2-4ac}}{2a} \\ \ \\ =\frac{6966 kg^2m/s ± \sqrt{(-6966 kg^2m/s)^2-4(795kg^2)(-109768 kg^2m^2/s^2))}}{2(795 kg^2)}\] giving two solutions \[v_2'=-8.16 m/s\] \[v_2'= 16.92 m/s\] We want the positive one, because the elephant was originally traveling in the positive direction, so the ball will be, too. Methinks the other way was better...........

OpenStudy (anonymous):

I'm sorry, this is way less than helpful :P